# Photometry of Saturated Stars with Machine Learning

### v1

@@ 5 Conclusions and Future Work
We developed a deep neural network (DNN) for the accurate photometry of saturated stars. The network was trained using predictions for the counts produced by the target star in a sample of $\sim 332$ thousand postage stamp images of stars roughly uniformly spread in magnitude from $g\sim 3$ to $15$ mag. For the verification sample there is little bias in the magnitude estimates except for the fainter stars, and 68% (95%) of the magnitude differences are less than $0.12$ (0.34)0.34)0.34 ) mag (see Fig. 3). More importantly, the typical scatter in the light curve of a saturated star is only $0.037$ mag (half the 15 to 85 percentile range about the mean) compared to $0.22$ mag for the SP2 light curves (see Fig. 4). The DNN light curves of many bright variable stars are dramatically better than their SP2 light curves (see Figs. 6 to 9) although the performance can some times be poor, as illustrated by the naked eye Cepheids in Fig. 7 (see below). Although the network was trained using data from only one camera, it appears to work equally well when tested using data from all cameras and both V and g band data when combined using a Gaussian process to intercalibrate the light curve means (see Figs. 10, 11). For some systems, this calibration procedure and/or outlier rejection would have to be done more carefully, but the results are otherwise very encouraging.
@@
Using the as-reduced images creates the problems seen for the bright Cepheids because of two issues. The first is that an interplay between adding bias corrections and the inheritance from the ASAS pipeline of reading only integer fits files can lead to damaging saturated images because of integer overflows. This manifests as pixels in the star being flagged as bad and given a pixel value of unity. This obviously causes severe problems for the standard aperture photometry pipeline. The DNN pipeline largely manages to correctly interpret the flagged pixels and recover a good estimate of the true flux. While we understand the origin of the problem, a fix is not trivial and has yet to be implemented. We believe the performance of a DNN retrained after the problem is fixed should be improved.
@@
One likely area where our results could be improved is in the estimation of the predicted counts for the training set. As a reminder, we predicted SDSS g magnitudes from Gaia DR3 magnitudes, converted them to counts using the zero point of the ASAS-SN reference image and the image subtraction estimate of the transparency difference between the images defining the zero point and the current image, and attempted to correct for the flux from other stars within the large ASAS-SN PSF. The Riello et al. (2021) transformation from Gaia magnitudes to g band is not great, with a reported scatter of $0.075$ mag. This is, however, not large enough to be a major contributer to the scatter seen for the verification data. Nonetheless, it might be better to have used ATLAS REFCAT (Tonry et al. 2018b) since it was built in part to systematically estimate $g$ band magnitudes.
The crowding corrections were computed based only on the Gaia G band magnitudes, and they should probably have been based on the estimated g band magnitudes. The thought was that these corrections should mostly be small, so modest fractional errors in them would not be important. Retrospectively, this may be need to be treated more carefully for the non-saturated stars. For the saturated stars which are the primary target of the project, this is less of an issue simply because the stars are so bright and the standard pipeline performs so poorly. Moreover, the scatter seen for the unsaturated stars in the verification data does not seem to depend on Galactic latitude, which suggests that problems with confusion cannot be a dominant driver of the scatter.
@@
Probably the most important issues to be addressed are the problems created by the ASAS-SN pipeline, fixing the damage done to saturated stars and either understanding the failures of the saturated star corrections or switching to using images without the corrections. We need to more extensively test how well the DNN works on multiple cameras to determine if we can simply train a single DNN for all cameras, perhaps with the camera identification as an additional item of metadata.
This method is now available as an option on ASAS-SN Sky Patrol v1.0 (https://asas-sn.osu.edu/, Kochanek et al. 2017) as the “Saturated Stars (Machine Learning)” photometry option. As should be clear from the examples, the results become less predictable for stars approaching naked eye visibility and brighter. The present method supplies no magnitude uncertainties, so these are simply reported as the typical scatter seen in Fig. 4. Better error estimates of the relative errors can be derived from the light curves, but this requires some information on the nature of any variability (e.g. periodic or non-periodic), making it difficult to estimate automatically. We anticipate that the machine learning results will steadily improve, particularly once the reduction pipeline problems for saturated stars are corrected. It is also considerably faster than the most equivalent mode of Sky Patrol v1.0 (“Aperture Photometry”), which takes roughly 5 times longer to produce the light curves shown in Figs. 6-9.

### v2

@@ 5 Conclusions and Future Work
We developed a multilevel perceptron (MLP) neural network for the accurate photometry of saturated stars. The network was trained using predictions for the counts produced by the target star in a sample of $\sim 332$ thousand postage stamp images of stars roughly uniformly spread in magnitude from $g\sim 3$ to $15$ mag. For the verification sample, the median magnitude differences are well within the 68% scatter of the differences ($0.12$ mag) except for the very brightest and faintest stars (see Fig. 3). More importantly, the typical scatter in the light curve of a saturated star is only $0.037$ mag (half the 15 to 85 percentile range about the mean) compared to $0.22$ mag for the SP2 light curves (see Fig. 4). The MLP light curves of many bright variable stars are dramatically better than their SP2 light curves (see Figs. 6 to 9) although the performance can sometimes be poor, as illustrated by the naked eye Cepheids in Fig. 7 (see below). Although the network was trained using data from only one camera, it appears to work equally well when tested using data from all cameras and both V and g band data when combined using a Gaussian process to intercalibrate the light curve means (see Figs. 10, 11). For some systems, this calibration procedure and/or outlier rejection would have to be done more carefully, but the results are otherwise very encouraging.
@@
Using the as-reduced images creates the problems seen for the bright Cepheids because of two issues. The first is that an interplay between adding bias corrections and the inheritance from the ASAS pipeline of reading only integer fits files can lead to damaging saturated images because of integer overflows. This manifests as pixels in the star being flagged as bad and given a pixel value of unity. This obviously causes severe problems for the standard aperture photometry pipeline. The MLP pipeline largely manages to correctly interpret the flagged pixels and recover a good estimate of the true flux. While we understand the origin of the problem, a fix is not trivial and has yet to be implemented. We believe the performance of a MLP retrained after the problem is fixed should be improved.
@@
One likely area where our results could be improved is in the estimation of the predicted counts for the training set. As a reminder, we predicted SDSS g magnitudes from Gaia DR3 magnitudes, converted them to counts using the zero point of the ASAS-SN reference image and the image subtraction estimate of the transparency difference between the images defining the zero point and the current image, and attempted to correct for the flux from other stars within the large ASAS-SN PSF. The Riello et al. (2021) transformation from Gaia magnitudes to g band is not great, with a reported scatter of $0.075$ mag. This is, however, not large enough to be a major contributor to the scatter seen for the verification data. Nonetheless, it might be better to have used ATLAS REFCAT (Tonry et al. 2018b) since it was built in part to systematically estimate $g$ band magnitudes.
The crowding corrections were computed based only on the Gaia G band magnitudes, and they should probably have been based on the estimated g band magnitudes. The thought was that these corrections should mostly be small, so modest fractional errors in them would not be important. Retrospectively, this may be need to be treated more carefully for the non-saturated stars. There might also be issues from the interplay between the estimated crowding corrections and the size and geometry of the postage stamp images (Fig. 1). For the saturated stars, which are the primary target of the project, this is less of an issue simply because the stars are so bright and the standard pipeline performs so poorly. Moreover, the scatter seen for the unsaturated stars in the verification data does not depend on Galactic latitude, which suggests that problems with confusion cannot be a dominant driver of the scatter.
@@
The most important issues to be addressed are the problems created by the ASAS-SN pipeline, fixing the damage done to the images of saturated stars and either understanding the failures of the saturated star corrections or switching to using images without the corrections. We need to more extensively test how well the MLP works on multiple cameras to determine if we can simply train a single MLP for all cameras, perhaps with the camera identification as an additional item of metadata.
This method is now available as an option on ASAS-SN Sky Patrol v1.0 (https://asas-sn.osu.edu/, Kochanek et al. 2017) as the “Saturated Stars (Machine Learning)” photometry option. As should be clear from the examples, the results become less predictable for stars approaching naked eye visibility and brighter. The present method supplies no magnitude uncertainties, so these are simply reported as the typical scatter seen in Fig. 4. Better error estimates of the relative errors can be derived from the light curves, but this requires some information on the nature of any variability (e.g., periodic or non-periodic), making it difficult to estimate automatically. We anticipate that the machine learning results will steadily improve, particularly once the reduction pipeline problems for saturated stars are corrected. It is also considerably faster than the most equivalent mode of Sky Patrol v1.0 (“Aperture Photometry”), which takes roughly 5 times longer to produce the light curves shown in Figs. 6-9.